the lenghts of the legs of a right triangle are 7 and 24. Find the lenght of the hypotenuse
Answer by Nate(3500)

use the theorm we all know!
c^2=a^2+b^2 where c is the lenght of hypontenuse and the legs are a and b
c^2=(7)^2+(24)^2
c^2=49+576
c^2=625
the hypontenuse is 25 units long